Last fall, the Wall Street Journal published a cost analysis of the new iPhone XS Max.
In round figures, the product price is $1100 per phone, and the marginal cost per phone is $400.
What is the current margin (Lerner Index, (P - MC)/P) for iPhone XS Max?
If Apple uses optimal pricing for this phone, what is the implied elasticity of demand for the iPhone XS Max?
Answer #1

Current margin = (P-MC)/P

Current margin = (1100-400)/1100

Current margin = 63.64%

----

Implied elasticity of demand = -1/current margin = -1/63.64%

Implied elasticity of demand = -1.57 or 1.57 (negative)